I. A sales manager receives 6 calls on average between 9:30 a.m. and 10:30 a.m. on a weekday. Find the probability that: a) he will receive 2 or more calls between 9:30 a.m. and 10:30 a.m. on a certain weekday. b) he wiI1 receive exactly 2 calls between 9:30 a.m. and 9:40 a.m. on a certain weekday. c) during a 5 day working week, there will be exactly 3 days on which he will receive no calls between 9:30 a.m. and 9:40 a.m. II. Suppose an ordinary six-sided die is rolled repeatedly and the outcome is noted on each roll. What is the probability that the: a) third 6 occurs on the seventh roll? b) number of rolls until the first 6 occurs is at most 10? 


1
Expert's answer
2021-07-01T06:16:30-0400

(1) here,

"\lambda=6"


(a) "P(X\ge 2)=1-[P(X=0)+P(X=1)]"


"=1-[\dfrac{e^{-\lambda} \lambda^0}{0!}+\dfrac{e^{-\lambda} \lambda^1}{1!}]
\\[9pt]
 =1-[\dfrac{e^{-6} 6^0}{0!}+\dfrac{e^{-6} 6^1}{1!}]
\\[9pt]
 =1-0.01735=0.98234"


(b) "P(X=2)=\dfrac{e^{-\lambda } \lambda^2}{2!}=\dfrac{e^{-6}. 6^2}{2!}=0.0446"


(c) In 3 days Probability that exactly 3 days on which he do not recieve any call-

"=3\times \dfrac{e^{-\lambda} \lambda^0}{0!}=3\dfrac{e^{-6} 6^0}{0!}=0.007436"


(2)


(a) Probability that 3 six occur on the 7 roll-


="^6C_2(\dfrac{1}{6})^2(\dfrac{5}{6})^4\times \dfrac{1}{6}=0.0334"



(b) Probability that the number of rolls until the first 6 occurs is at most 10-


"=1-^{10}C_{0} (\dfrac{1}{6})^0(\dfrac{5}{6})^{(10-0)}\times \dfrac{1}{6}=1-0.026917=0.97308"